Explain in words what happens to the multiplier when an economy moves from a small closed economy with a government to an open economy. What is the formula for the multiplier in a small open economy?
Does the size of the multiplier increase or decrease?
Show the formula of the multiplier
1
Expert's answer
2020-06-08T11:08:07-0400

multiplier in a closed economy:

k= 1 / MPS

multiplier in a semi-open economy:

k= 1 / (MPS + MPM)

multiplier in a semi-closed economy:

k = / (MPS – МРХ)

multiplier in an open economy:

k= 1/(1-MPC) = 1 / (MPS МРМ – МРХ)


The value of the multiplier calculated for an open economy will be less than for a closed one. This is because the cost of importing products, as well as savings, cease to be an integral part of the aggregate demand for the domestic product.